Alex, Stuart and Ali recieve an average of $12 per week in pocket money. Alex recieves $2 more than Stuart, but $5 less than Ali. How much does each one recieve?
Could you please help me answer this question

Answer by rfer(16322)

Alex=x
Stuart=x-2
Ali=x+5
Average of 12x3=36
x+x-2+x+5=36
3x=33
x=11
Alex=11
Stuart=11-2=9
Ali=11+5=16